a balloon contains gas at a pressure 1.2 atm (1 atm=101.3kpa) and has a volume of 0.10m^3. more gas is pumped into the balloon at a constant pressure until the volume is doubled. how much work is done by the pump? A) 12 J B) 24 KJ C) 24 J D)12 KJ E) 6.1 KJ
Answer #1

we have realtion for work done W=PV

where P is the pressure and V is the change in the volume.

P=1.2101.3 =121.56 Kpa , V =(0.20 - 0.10 )=0.10 m3

W=(121.56)(0.10)

W=12.15 KJ

so answer is D 12 K J